OpenStudy (anonymous):

| x + 3 | >= 7 Solve and graph the inequality.

OpenStudy (anonymous):

\[|x + 3| \ge 7\] \[|x + 3| = \text{Everything between | and | is positive.}\] Once you made everything positive (remove the absolute value markers): \[x + 3 \ge 7 \] Solve for \(x\): \[x + 3 - 3 \ge 7 - 3\] \[x \ge 4 \] So \(x\) is greater than or equal to 4.